OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y less than or greater to x - 5 y greater than or equal to -x - 4 (-5, 2) (5, -2) (-5, -2) (5, 2)

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

please someone can help me

OpenStudy (anonymous):

(5,-2) your solution is on the region under the lin Y=x-5 and over the line y=-x-4